OpenStudy (anonymous):

Solve using the elimination method. Show your work. If the system has no solution or an infinite number of solutions, state this. 4x + 10y = -1 8x – 6y = -15

OpenStudy (anonymous):

-8x(4x +10y)= -1 4x(8x - 6y)= -15 -32x-80y = 8 +32x-24y= -60 -104y = -52 -104y/-104y = -52/-104 Y = .5 or 1/2 Substitute .5 into first original equation. 4x + 10(.5) = -1 4x + 5 = -1 4x +5 -5 = -1 - 5 4x = -6 4x/4 = -6/4 x = 1.5

OpenStudy (anonymous):

is this right

OpenStudy (anonymous):

x = -1.5, NOT 1.5 Y is correct

OpenStudy (anonymous):

pk51 is correct

OpenStudy (anonymous):

ty